### v1

# Argument Rarity-based Originality Assessment for AI-Assisted Writing

Abstract
As Large Language Models (LLMs) have become capable of effortlessly generating high-quality text, traditional quality-focused writing assessment is losing its significance. If the essential goal of education is to foster critical thinking and original perspectives, assessment must also shift its paradigm from quality to originality. This study proposes Argument Rarity-based Originality Assessment (AROA), a framework for automatically evaluating argumentative originality in student essays. AROA defines originality as rarity within a reference corpus and evaluates it through four complementary components: structural rarity, claim rarity, evidence rarity, and cognitive depth. The framework quantifies the rarity of each component using density estimation and integrates them with a quality adjustment mechanism, thereby treating quality and originality as independent evaluation axes. Experiments using human essays and AI-generated essays revealed a strong negative correlation between quality and claim rarity, demonstrating a quality-originality trade-off where higher-quality texts tend to rely on typical claim patterns. Furthermore, while AI essays achieved comparable levels of structural complexity to human essays, their claim rarity was substantially lower than that of humans, indicating that LLMs can reproduce the form of argumentation but have limitations in the originality of content.
keywords:
Argumentation Mining , Large Language Models , Automated Essay Scoring , Writing Assessment , AI in Education ,
journal: Computers Education: Artificial Intelligence
[…]
[1]organization=Ritsumeikan Global Innovation Research Organization (R-GIRO), Ritsumeikan University, addressline=1-1-1, Nojihigashi, city=Kusatsu, postcode=525-8577, state=Shiga, country=Japan
[…]
[2]organization=Faculty of Business and Commerce, Kansai University, addressline=3-3-35, Yamatecho, city=Suita, postcode=564-8680, state=Osaka, country=Japan
[…]
[3]organization=Arbege Corporation, addressline=3-20, Yotsuyatori, city=Nagoya, postcode=464-0819, state=Aichi, country=Japan
[…]
[4]organization=College of Life Sciences, Ritsumeikan University, addressline=1-1-1, Nojihigashi, city=Kusatsu, postcode=525-8577, state=Shiga, country=Japan
[…]
[5]organization=Office of General Education, Ritsumeikan University, addressline=1-1-1, Nojihigashi, city=Kusatsu, postcode=525-8577, state=Shiga, country=Japan
1 Introduction
The rapid advancement of Large Language Models (LLMs) is fundamentally transforming the nature of human intellectual activities [8, 35]. LLMs, exemplified by ChatGPT, have acquired the ability to instantly generate logically coherent, fluent, and persuasive text on any given topic. The capability to write well, once the exclusive domain of humans, is no longer unique to human beings. While this technological innovation is affecting various domains of society, its impact is particularly profound in educational settings. Determining whether an essay or report submitted by a student is truly the product of original thinking or merely a reproduction of typical patterns has become more challenging than ever before.
Looking back at the history of writing assessment, evaluation criteria have consistently emphasized quality. Logical coherence, grammatical accuracy, clarity of organization, and persuasiveness of argumentation have been widely recognized as universal characteristics of excellent writing. Research on Automated Essay Scoring (AES) has primarily focused on predicting these quality indicators [18]. However, now that LLMs have reached a level where they surpass humans in quality, assessment based solely on such criteria is entering a stage where its fundamental significance must be reconsidered. If AI can generate essays of near-perfect quality, high quality alone is no longer a sufficient indicator reflecting human thinking ability or learning outcomes.
What should be questioned here is the essential purpose of having students write in education. The true purpose of assigning essays to students is not the production of well-formed text itself. Rather, it lies in having students experience the process of critical thinking: thinking deeply about a given topic, forming original perspectives, exploring diverse evidence, and anticipating and responding to counterarguments [13]. In other words, what education should originally evaluate is not the quality of writing but the originality of the thinking behind it. From this perspective, the paradigm of assessment must shift from quality to originality.
What then does argumentative originality specifically refer to? This study conceptualizes originality not simply as being different from others in a superficial sense, but as multifaceted characteristics of argumentation. First, there is the uniqueness of the argument structure itself. This dimension concerns whether the structural pattern of how claims are supported by reasons, how evidence is arranged, and how counterarguments are addressed is typical or rare. Second, there is the uniqueness of claim content. This dimension concerns whether the claim is one that many people would make on the same topic or whether it represents an original perspective. Third, there is the uniqueness of evidence selection. This dimension concerns whether the argument relies on commonly cited typical evidence or employs rare evidence that was independently explored. Fourth, there is the depth of reasoning. This dimension concerns whether the argument remains at a simple claim-evidence correspondence or develops multi-layered reasoning that anticipates counterarguments and constructs rebuttals. These dimensions are mutually independent, and being original in one dimension does not guarantee originality in others. For truly comprehensive originality assessment, these multifaceted aspects must be captured individually and integrated.
However, consistent evaluation of such argumentative originality by human raters involves inherent difficulties. In educational settings, instructors must grade dozens to hundreds of essays within limited time at the end of a semester. Under such circumstances, it is difficult to avoid order effects and fatigue in grading, where impressions from earlier essays unconsciously influence subsequent evaluations [19]. Additionally, the evaluator’s own expertise and prior knowledge act as biases [1]. Evaluators tend to give higher scores to claims that align with their expectations and beliefs, while finding it difficult to appropriately recognize the value of unfamiliar perspectives or rare evidence. As an even more fundamental problem, judging how rare an essay’s claims or evidence are within the entire corpus when evaluating individual essays exceeds the limits of human cognitive capacity. It is practically impossible for evaluators to simultaneously remember the content of hundreds of essays and judge relative rarity. Due to these limitations, originality assessment tends to be subjective and inconsistent, making it difficult to provide fair and meaningful feedback to students.
In recent years, new possibilities have emerged through leveraging the advanced language understanding capabilities of LLMs. LLMs possess the ability to deeply understand the semantic content of text, analyze argument structure, and grasp relationships between elements [20]. By applying this capability to assessment, there is potential to complement the limitations of human evaluators. LLMs do not fatigue, are not affected by order effects, and can evaluate all essays according to the same criteria. They can also process large-scale corpora and objectively calculate the relative positioning of each essay [3]. Indeed, such AI-assisted assessment is already being implemented in practice. For example, at Ritsumeikan University, a large private university in Japan, on-demand data science courses accommodate up to 1,000 students per class. To address this scale, the authors have independently developed and deployed an AI-powered assessment system called RaaS (Ritsumeikan AI-powered Assessment Solution), which is already operational in actual courses[25]. While final evaluation responsibility remains with human instructors, this represents an emerging case of large-scale course operation premised on AI-assisted assessment. Despite such practical implementations, most existing research on AI-assisted assessment has focused on quality evaluation and grammatical error detection, and no framework has been established for systematically evaluating originality. The framework for quality assessment cannot be directly applied to originality assessment. Originality is an essentially different concept from quality and requires its own theoretical foundation and evaluation methodology.
This study proposes Argument Rarity-based Originality Assessment (AROA), a framework for automatically evaluating argumentative originality in argumentative essays. The core idea of AROA is to operationally define originality as rarity within a reference corpus. For a claim in an essay to be original means that the claim is rare among other essays addressing the same topic. This definition makes it possible to measure originality, an abstract concept, as a quantitative indicator based on density estimation.
AROA adopts a multifaceted approach that conceptualizes originality as four complementary components. The first component, structural rarity, quantifies how rare the pattern of argument structure consisting of claims, reasons, evidence, counterarguments, and rebuttals is within the reference corpus, using kernel density estimation. The second component, claim rarity, measures how far the central claims of an essay are from other essays in semantic space, using neighborhood density of sentence embeddings. The third component, evidence rarity, evaluates the uniqueness of evidence used to support claims using the same methodology. The fourth component, cognitive depth, quantifies the dialogic structure of counterargument-rebuttal exchanges and the hierarchical depth of reasoning as complexity of cognitive operations. These four components have been confirmed through correlation analysis to have low mutual correlations, each capturing different aspects of originality. AROA integrates these components after adjusting them with quality scores, thereby treating quality and originality as independent axes while calculating a comprehensive originality score that considers the balance between both.
The contributions of this study are summarized in the following three points.
  1. (i)
    AROA is a novel framework designed to automatically evaluate argumentative originality in argumentative essays through four complementary components—structural rarity, claim rarity, evidence rarity, and cognitive depth—integrating density-based rarity quantification with quality adjustment.
  2. (ii)
    Empirical demonstration of a strong negative correlation between text quality and semantic rarity through large-scale experiments using human essays and AI-generated essays, revealing a quality-originality trade-off where logically coherent, high-quality texts tend to rely on typical claim patterns and common evidence.
  3. (iii)
    Comparative analysis of component profiles between human essays and AI essays, revealing that contemporary LLMs achieve comparable levels to humans in structural complexity of argumentation while exhibiting notable limitations in semantic originality of claims and evidence, thereby presenting new guidelines for originality assessment in the AI era.
The remainder of this paper is organized as follows. Section 2 reviews related work and clarifies the positioning of this study. Section 3 describes the details of the proposed framework AROA. Section 4 presents the experiments and results, and Section 5 discusses the implications and limitations of the results. Finally, Section 6 presents the conclusions.
2 Related Work
2.1 Automated Essay Scoring
AES is a technology that automatically assigns scores to essays using computers. From the late 1990s to the 2000s, practical systems such as e-rater [7], IntelliMetric, and Intelligent Essay Assessor were developed and deployed in large-scale standardized tests such as GMAT and GRE. Early AES systems extracted handcrafted features such as lexical diversity, syntactic complexity, and discourse structure, and predicted scores using regression models. However, Powers et al. [24] conducted experiments in which professional writers intentionally created essays to deceive e-rater, demonstrating that texts that were content-poor but satisfied formal features tended to receive high scores. This study made it clear that AES at the time relied heavily on surface features and could not adequately evaluate the validity of logic or depth of content.
Since 2016, with the advancement of deep learning, the technical paradigm of AES has undergone a major transformation. Taghipour and Ng [30] proposed an end-to-end AES using LSTM and demonstrated the effectiveness of neural approaches that eliminate the need for handcrafted feature engineering. Hussein et al. [15] provided a comprehensive review that systematically classified AES research into handcrafted feature-based and automatic feature learning-based approaches, identifying three major challenges of AES: human-likeness, deception resistance, and creativity assessment.
In recent years, the application of large language models (LLMs) such as GPT to AES has rapidly progressed. Naismith et al. [20] proposed a method to evaluate discourse coherence in essays using GPT-4 and demonstrated the possibility of explainable AES that can also generate evaluation rationales. Xiao et al. [36] demonstrated that novice evaluators who referred to LLM feedback achieved accuracy close to that of experts, presenting a perspective that redefines AES not as complete automation but as human capability augmentation. However, Bevilacqua et al. [4] reported that LLM-based AES tends to rate GPT-generated texts higher than human texts, clarifying a fundamental bias problem of AES in the generative AI era.
Surveying these prior studies reveals that AES research has consistently focused on predicting text quality—logical coherence, grammatical accuracy, and clarity of organization. From early handcrafted feature-based systems to the latest LLM-based approaches, the evaluation target has been whether text is well-written, not whether argumentation is original. As Dikli [10] pointed out, creativity and critical thinking assessment have been recognized as challenges for AES, yet no framework has been proposed to systematically address this. AROA, proposed in this study, addresses this research gap by introducing argumentative originality as an evaluation axis independent of quality, presenting a new direction for AES research.
2.2 Argumentation Mining
Argumentation Mining is a research field that automatically identifies and extracts argument structures from natural language text. Lawrence and Reed [16] provided a comprehensive survey of this field, systematically organizing topics from the identification of claims and premises to the extraction of support and rebuttal relations, and the mapping to the Toulmin model and argumentation schemes. The argumentation model proposed by Toulmin [31] is widely adopted as the theoretical foundation of argumentation mining. This model structures argumentation through six elements—Claim, Grounds (Data), Warrant, Backing, Qualifier, and Rebuttal—and has become the standard framework for argumentative essay analysis in the educational domain.
Argumentation mining research targeting student essays was formally initiated by Stab and Gurevych [29]. They proposed an annotation scheme for claims, premises, and support/attack relations in persuasive essays, achieving high inter-annotator agreement. This study became the starting point for subsequent research on argument structure analysis in student essays. Wachsmuth et al. [33] proposed a method to evaluate argument quality (organization, strength) in essays using extracted argument structures, extending argumentation mining from structure extraction to quality assessment.
Educational applications of argumentation mining have also progressed. Nguyen and Litman [21] integrated argument features such as the number of claims and support relations into AES and demonstrated improved accuracy in persuasive essay scoring. Wambsganss et al. [34] developed an educational support system that visualizes argument structures and provides feedback to students, showing that pointing out missing claims or insufficient evidence promotes learners’ argumentative self-correction.
In recent years, argumentation mining research using LLMs has rapidly progressed. Naismith et al. [20] proposed a method to evaluate discourse coherence in essays using GPT-4 and demonstrated that it can better reflect argument structure quality than conventional surface cohesion indicators. Li et al. [17] proposed a multi-task benchmark integrating claim detection, evidence extraction, and persuasiveness evaluation, revealing that even LLMs with high performance on single tasks show significant performance degradation when performing argument understanding in a chained manner.
These prior studies have made important advances in argument structure extraction and quality assessment. However, the quality evaluated in existing research concerns aspects such as structural completeness, organization, and persuasiveness, and does not include the originality of argument content, that is, how rare claims and evidence are compared to other essays. AROA, proposed in this study, utilizes argumentation mining techniques to extract claims, evidence, counterclaims, and rebuttals, and then quantifies their semantic rarity through density estimation. This presents a new direction of evaluating originality rather than quality of argument structures.
2.3 Originality and Creativity Assessment
Originality and creativity assessment has a long tradition of research in psychology and education. Runco and Yoruk [27] defined creativity as the combination of novelty and usefulness, and demonstrated that divergent thinking is the representative paradigm for measuring this. Among divergent thinking tests, the Alternative Uses Test (AUT) is the most widely used, asking respondents to generate as many alternative uses as possible for a given object, and evaluating responses in terms of fluency, flexibility, originality, and elaboration.
Traditionally, originality assessment in AUT was conducted by human raters subjectively judging the rarity of responses, but in recent years, automatic evaluation based on semantic distance has become mainstream. Beaty et al. [2] proposed a method to automatically evaluate originality by calculating the semantic distance between stimulus words and response words using sentence embeddings. This method showed moderate to high correlation with human evaluation and was established as the standard approach for automatic creativity assessment. Patterson et al. [23] verified the multilingual validity of semantic distance-based originality assessment using large-scale data from 12 languages and over 6,500 participants, demonstrating that Transformer-based embeddings enable stable evaluation across languages.
In recent years, creativity assessment using LLMs has rapidly developed. DiStefano et al. [9] fine-tuned LLMs for the higher-order creative language task of metaphor generation and achieved agreement with human evaluation that significantly exceeded semantic distance methods. Valueva et al. [32] comprehensively reviewed the evolution of automatic scoring methods for divergent thinking tests, pointing out that while LLMs are approaching human evaluation, the opacity of judgment rationales is a new challenge. On the other hand, Runco et al. [26] analyzed divergent thinking responses generated by GPT-3.5 and GPT-4, showing that while AI can generate numerous ideas, it also includes many irrelevant and redundant expressions. This result reconfirms the importance of the definition that creativity equals novelty plus usefulness.
These prior studies have succeeded in evaluating the novelty of short responses such as those in divergent thinking tests. However, no research exists that evaluates argumentative originality in long texts such as argumentative essays, that is, how rare argument elements such as claims, evidence, and counterarguments are. Existing creativity assessment research measures semantic distance of words or short sentences, and no framework has been proposed to capture the originality of entire argument structures in a multifaceted manner. AROA, proposed in this study, integrates the concept of rarity-based evaluation with argumentation mining, assessing argumentative originality through four components: claim rarity, evidence rarity, structural rarity, and cognitive depth. This presents a new direction that extends the concept of rarity as originality, established in divergent thinking research, to complex texts such as argumentative essays.
3 Argument Rarity-based Originality Assessment
3.1 Framework Overview
This study proposes AROA, a structure-oriented originality assessment framework that evaluates essays based on the originality of achieved thinking rather than surface fluency or author attribution, in an environment where AI-assisted essay writing has become commonplace. Fig. 1 shows the overall processing pipeline of AROA.
Figure 1: Overview of the AROA processing pipeline.
The design philosophy of AROA is based on Toulmin’s [31] argumentation structure model and integrates insights from critical thinking education [11, 12] and deep learning theory [5]. In particular, in the current situation where AI-generated text surpasses humans in surface fluency, the necessity of structurally evaluating learners’ cognitive depth is also supported from the perspective of knowledge construction theory [28].
The AROA processing pipeline consists of five stages. First, given a collection of essays on the same topic $\mathcal{D}=\{x_{i}\}_{i=1}^{N}$, each essay is transformed using an LLM into a semantic structure $\mathcal{S}(x)$ consisting of five elements: Claim, Reason, Evidence, Counterargument, and Rebuttal. Next, each extracted element is converted into a $d$-dimensional embedding vector using Sentence-BERT, enabling quantitative computation of semantic similarity between elements. Subsequently, based on the embedding representations and structural information, originality sub-scores are calculated from the following four perspectives:
  1. i)
    Structural rarity ($S_{\text{struct}}$): Evaluates the atypicality of reasoning structure through density estimation
  2. ii)
    Claim rarity ($S_{\text{claim}}$): Evaluates the semantic rarity of central claims using the $K$-nearest neighbor method
  3. iii)
    Evidence rarity ($S_{\text{evid}}$): Evaluates the use of specific and rare evidence
  4. iv)
    Cognitive depth ($S_{\text{move}}$): Evaluates the correspondence between counterarguments and rebuttals, and the depth of reasoning paths
Furthermore, to prevent essays that are original but logically flawed from receiving high scores, a quality score $Q(x)$ that does not depend on fluency is calculated. Finally, the four originality sub-scores are z-score normalized and then integrated through weighted combination, multiplied by the quality score to compute the final evaluation value $S_{\text{final}}(x)$.
The rationale for selecting these four perspectives is based on the fact that originality is a multifaceted concept that cannot be captured by a single dimension. Structural rarity evaluates the uniqueness of how arguments are made, claim and evidence rarity evaluate the uniqueness of what is argued, and cognitive depth evaluates the uniqueness of how deeply arguments are developed. This multifaceted approach enables evaluation of originality accompanied by quality as argumentation, rather than superficial novelty. The following sections explain the details of each processing stage in order.
3.2 Argument Structure Extraction
This section explains the process of extracting argument structure from essays and converting it into a format suitable for subsequent rarity assessment. First, for each essay $x$, we define a structural representation consisting of the following five elements:
$\mathcal{S}(x)=\bigl\{C(x),\;R(x),\;E(x),\;A(x),\;B(x)\bigr\}$ (1)
The definition of each element is shown in Table 1. This five-element representation is an adaptation of Toulmin’s [31] argumentation structure model to the context of educational assessment. While Toulmin’s original model defines six elements, Claim, Data, Warrant, Backing, Qualifier, and Rebuttal, this study adopts a five-element model that integrates Warrant and Backing into Reasons and omits Qualifier. This simplification aims to improve practicality in educational settings and consistency of LLM extraction.
Table 1: Definition of five argument elements
Symbol Name Description
$C(x)=\{c_{k}\}_{k=1}^{n_{C}}$ Claims Central conclusions or assertions that the author wants the reader to accept
$R(x)=\{r_{k}\}_{k=1}^{n_{R}}$ Reasons Arguments that support the claims; fundamental grounds for why the claims should be accepted
$E(x)=\{e_{k}\}_{k=1}^{n_{E}}$ Evidences Examples, data, statistics, or citations that back up the reasons
$A(x)=\{a_{k}\}_{k=1}^{n_{A}}$ Counterarguments Anticipated opposing views, criticisms, or alternative interpretations
$B(x)=\{b_{k}\}_{k=1}^{n_{B}}$ Rebuttals Responses to counterarguments; refutations or re-assertions
These five elements form hierarchical support relationships. Reasons support Claims, and Evidences back up Reasons. Additionally, Counterarguments oppose Claims, and Rebuttals respond to Counterarguments. This relational structure enables capturing not only the presence or absence of elements but also the logical composition of argumentation.
An LLM is used for structure extraction. The input to the LLM consists of a two-layer structure: a system prompt and a user prompt. In the system prompt, the definition and guidelines for the extraction task were specified as follows:
You are an expert in argument structure analysis. Extract the following five elements from the given essay: (1) Claims: Central conclusions or assertions of the essay, (2) Reasons: Arguments that support the claims, (3) Evidences: Examples, data, or specific information that back up the reasons, (4) Counterarguments: Anticipated opposing views or alternative interpretations, (5) Rebuttals: Responses to counterarguments, re-assertions. Also identify relationships between elements: supports (reason supports claim, evidence backs reason), opposes (counterargument opposes claim), responds_to (rebuttal responds to counterargument). Guidelines: Preserve the original meaning; avoid excessive summarization. Split compound sentences into semantic units. If an element does not exist, use an empty array.
The three guidelines specified aim to ensure that extracted elements accurately reflect the intent of the original text, to unify element granularity by splitting compound sentences into semantic units, and to allow empty sets when elements do not exist. In the user prompt, the essay text was provided as input with instructions for structured output in JSON format. The output format includes text lists of each element as well as relational information between elements. To stabilize LLM output, the temperature parameter was set to $T=0$ to obtain deterministic output.
From the extracted structure, two derived features used in subsequent evaluation are calculated. First, the depth of reasoning $\text{Depth}(x)$ represents the hierarchical complexity of argumentation and is defined as follows:
[…]
This definition progressively distinguishes from simple argumentation with only claims (depth 1) to complex argumentation with evidence backing or counterargument-rebuttal dialogic structure (depth 3).
Second, the number of counterargument-rebuttal matched pairs $n_{\text{matched}}$ is counted as the number of responds_to relationships extracted. When relationships are not explicitly extracted, it is estimated as $n_{\text{matched}}=\min(n_{A},n_{B})$. This feature captures the completeness of thinking, whether counterarguments are not only presented but also appropriately responded to.
Each obtained text element is projected into vector space by the sentence embedding function $f:\text{Text}\to\mathbb{R}^{d}$:
$\mathbf{v}(c_{k}),\;\mathbf{v}(r_{k}),\;\mathbf{v}(e_{k}),\;\mathbf{v}(a_{k}),\;\mathbf{v}(b_{k})\in\mathbb{R}^{d}$ (3)
This study used a pre-trained Sentence-BERT model [14] to obtain $d=768$-dimensional embeddings. Sentence-BERT has been widely validated for its performance in capturing sentence-level semantic similarity and is suitable for this task. These embedding representations are used as the computational basis for semantic similarity in the claim and evidence rarity assessments explained in subsequent sections.
3.3 Structural Rarity Assessment
This section explains the method for calculating the structural rarity score $S_{\text{struct}}$, which quantifies how rare an essay’s argument structure is within the corpus. First, a 12-dimensional structural feature vector $\boldsymbol{\phi}(x)$ is constructed from each essay $x$:
$\boldsymbol{\phi}(x)=\begin{pmatrix}n_{C}\\ n_{R}\\ n_{E}\\ n_{A}\\ n_{B}\\ n_{R}/\max(n_{C},1)\\ n_{E}/\max(n_{R},1)\\ \text{Depth}(x)\\ \text{Branches}(x)\\ \mathbb{I}_{n_{A}>0}\\ \mathbb{I}_{n_{B}>0}\\ n_{\text{matched}}\end{pmatrix}$ (4)
Here, $n_{C},n_{R},n_{E},n_{A},n_{B}$ are the counts of each element, $n_{R}/\max(n_{C},1)$ is the number of reasons per claim, and $n_{E}/\max(n_{R},1)$ is the number of evidences per reason. $\text{Depth}(x)$ and $n_{\text{matched}}$ are the depth of reasoning and the number of counterargument-rebuttal matched pairs defined in the previous section. $\text{Branches}(x)=\max(n_{C},1)\times\max(n_{R},1)$ represents the number of argument branches, capturing multifaceted argument structures that support multiple claims with multiple reasons. $\mathbb{I}_{n_{A}>0}$ and $\mathbb{I}_{n_{B}>0}$ are indicator functions showing the presence or absence of counterarguments and rebuttals. These features are designed to capture multiple aspects: quantitative aspects of argumentation (element counts), structural aspects (ratios, depth, branches), and critical thinking aspects (existence and correspondence of counterargument-rebuttal).
The structural rarity score is defined as the negative logarithm of probability density in the feature vector space. This definition corresponds to self-information in information theory, where essays located in low-density regions receive higher scores. Prior to density estimation, to eliminate scale differences between features, standardization is performed on the set of feature vectors across the entire corpus $\Phi=\{\boldsymbol{\phi}(x_{i})\}_{i=1}^{N}$:
$\boldsymbol{\phi}^{\prime}(x)=\frac{\boldsymbol{\phi}(x)-\boldsymbol{\mu}_{\Phi}}{\boldsymbol{\sigma}_{\Phi}}$ (5)
Here, $\boldsymbol{\mu}_{\Phi}$ and $\boldsymbol{\sigma}_{\Phi}$ are the corpus-wide mean and standard deviation of each feature. Next, to mitigate the curse of dimensionality associated with density estimation in high-dimensional spaces, dimensionality is reduced to 3 using Principal Component Analysis (PCA). Kernel Density Estimation (KDE) is used for density estimation, with bandwidth automatically set based on Scott’s rule. The final structural rarity score is defined as follows:
$S_{\text{struct}}(x)=-\log\hat{p}(\boldsymbol{\phi}^{\prime}(x))$ (6)
Here, $\hat{p}(\boldsymbol{\phi}^{\prime}(x))$ is the KDE density estimate for the feature vector after standardization and PCA. This score causes essays with typical argument structures within the corpus to receive low scores, while essays with rare argument structures receive high scores.
3.4 Semantic Rarity Assessment
This section explains the method for calculating the semantic rarity scores $S_{\text{claim}}$ and $S_{\text{evid}}$, which quantify how rare the content of claims and evidence are within the corpus. While the structural rarity in the previous section evaluates how arguments are made, the semantic rarity in this section evaluates what is argued.
3.4.1 Claim Rarity
Claim rarity is evaluated based on how isolated each claim’s embedding vector is in semantic space. Let the set of embeddings for all claims in the corpus be $\mathcal{V}_{C}=\{\mathbf{v}(c)\mid c\in\bigcup_{i}C(x_{i})\}$. For each claim $c_{k}$, local density is calculated using the $K$-nearest neighbor method. First, the $K$ nearest neighbors of $\mathbf{v}(c_{k})$ are found, and local density is defined as the average cosine similarity with them:
$\text{density}(\mathbf{v}(c_{k}))=\frac{1}{K}\sum_{j=1}^{K}\text{sim}(\mathbf{v}(c_{k}),\mathbf{v}^{(j)})$ (7)
Here, $\mathbf{v}^{(j)}$ is the $j$-th nearest neighbor of $\mathbf{v}(c_{k})$, and $\text{sim}(\cdot,\cdot)$ is cosine similarity:
$\text{sim}(\mathbf{u},\mathbf{v})=\frac{\mathbf{u}\cdot\mathbf{v}}{\|\mathbf{u}\|\|\mathbf{v}\|}$ (8)
The claim rarity score is defined as the complement of density:
$\text{Rar}_{C}(c_{k})=1-\text{density}(\mathbf{v}(c_{k}))$ (9)
This design results in low scores when there are many similar claims nearby and high scores when there are few similar claims.
The claim rarity score for essay $x$ is calculated as the average rarity of all claims contained in that essay:
$S_{\text{claim}}(x)=\frac{1}{|C(x)|}\sum_{c_{k}\in C(x)}\text{Rar}_{C}(c_{k})$ (10)
When no claims exist ($|C(x)|=0$), $S_{\text{claim}}(x)=0$. This study used $K=5$.
3.4.2 Evidence Rarity
Evidence rarity is calculated using the same method as claims. Let the set of embeddings for all evidence in the corpus be $\mathcal{V}_{E}=\{\mathbf{v}(e)\mid e\in\bigcup_{i}E(x_{i})\}$, and define the rarity of each evidence $e_{k}$ as follows:
$\text{Rar}_{E}(e_{k})=1-\text{density}(\mathbf{v}(e_{k}))$ (11)
The evidence rarity score for essay $x$ is calculated as the average rarity of all evidence:
$S_{\text{evid}}(x)=\frac{1}{|E(x)|}\sum_{e_{k}\in E(x)}\text{Rar}_{E}(e_{k})$ (12)
When no evidence exists ($|E(x)|=0$), $S_{\text{evid}}(x)=0$.
These semantic rarity scores result in low scores for essays that use typical claims or evidence that frequently appear within the corpus, and high scores for essays that use original perspectives or rare evidence.
3.5 Cognitive Depth Assessment
This section explains the method for calculating the cognitive depth score $S_{\text{move}}$, which evaluates the depth of reasoning and completeness of response to counterarguments in argumentation. While the rarity assessments in previous sections measure whether something is rare, the cognitive depth in this section evaluates how deeply arguments are developed.
The cognitive depth score consists of three components:
$S_{\text{move}}(x)=\alpha\cdot n_{\text{matched}}+\beta\cdot\text{Depth}(x)+\gamma\cdot\text{Coverage}(x)$ (13)
Here, $\alpha,\beta,\gamma$ are weight parameters for each component. These three components capture different aspects of critical thinking: $n_{\text{matched}}$ evaluates the quantity of critical dialogue, $\text{Depth}(x)$ evaluates the hierarchical complexity of argumentation, and $\text{Coverage}(x)$ evaluates the completeness of critical dialogue. Each component is explained below.
The first component, $n_{\text{matched}}$, is the number of pairs where counterarguments and rebuttals semantically correspond. As explained in Section 3.2, this value is extracted as responds_to relationships during structure extraction, or estimated as $n_{\text{matched}}=\min(n_{A},n_{B})$ when relationships are not explicitly extracted. Therefore, a higher number of counterargument-rebuttal pairs indicates more active back-and-forth of critical thinking.
The second component, $\text{Depth}(x)$, is the depth of reasoning defined in Section 3.2, with 1 for claims only, 2 when reasons exist, and 3 when evidence exists or both counterarguments and rebuttals exist. This component evaluates hierarchical argument structures where claims are supported by reasons and further backed by evidence.
The third component, $\text{Coverage}(x)$, is the counterargument coverage rate, representing the proportion of presented counterarguments that are responded to with rebuttals:
$\text{Coverage}(x)=\begin{cases}\frac{n_{\text{matched}}}{n_{A}}&\text{if }n_{A}>0\\ 0&\text{otherwise}\end{cases}$ (14)
This metric captures the completeness of thinking, i.e., whether counterarguments are not only raised but also appropriately responded to. Essays that provide rebuttals to raised counterarguments rather than leaving them unaddressed receive higher scores.
This study used $\alpha=\beta=\gamma=1.0$. These weights represent a design that treats the three components equally, but they can be adjusted according to educational objectives. For example, if critical thinking is emphasized, $\alpha$ and $\gamma$ can be increased, and if hierarchical argumentation is emphasized, $\beta$ can be increased.
3.6 Quality Adjustment
This section explains the method for calculating the quality score $Q(x)$, which prevents essays that are original but logically flawed from receiving high scores. Since essays that are logically incoherent cannot be called original even if they have high rarity, adjustment by quality is necessary. The quality score consists of two components: Coherence and Logical Validity:
$Q(x)=w_{\text{coh}}\cdot Q_{\text{coherence}}(x)+w_{\text{log}}\cdot Q_{\text{logical}}(x)$ (15)
Here, $w_{\text{coh}}$ and $w_{\text{log}}$ are the weights for each component. This study used $w_{\text{coh}}=w_{\text{log}}=0.5$.
As an important design decision, this framework intentionally excludes Fluency from quality evaluation. The reason is that text generated by LLMs generally has high fluency, so including fluency in the evaluation would favor AI-generated text, contradicting the objective of this framework to evaluate originality of thinking. Each component is evaluated using an LLM. For coherence evaluation, judgment is made based on three criteria: i) whether there are contradictions between claims and reasons, ii) whether there are contradictions between reasons and evidence, and iii) whether there are contradictions among multiple claims. Output is in three levels: 1.0 (no contradictions), 0.5 (partial contradictions), and 0.0 (serious contradictions). For logical validity evaluation, judgment is made based on three criteria: i) whether reasons logically support claims, ii) whether evidence appropriately backs up reasons, and iii) whether there are logical leaps or fallacies. Output is in three levels: 1.0 (valid), 0.5 (partial problems), and 0.0 (serious defects).
In both evaluations, the prompt explicitly instructs not to evaluate fluency or style, ensuring that evaluation focuses only on logical structure. This evaluates only the quality as argumentation rather than the superficial smoothness of text.
3.7 Final Score Computation
This section explains the method for integrating the four originality sub-scores and quality score to compute the final evaluation value $S_{\text{final}}(x)$. The four sub-scores ($S_{\text{struct}}$, $S_{\text{claim}}$, $S_{\text{evid}}$, $S_{\text{move}}$) each have different scales, making direct comparison and integration difficult. To solve this problem, each sub-score is z-score normalized within the corpus:
$\tilde{S}_{\text{comp}}(x)=\frac{S_{\text{comp}}(x)-\mu_{\text{comp}}}{\sigma_{\text{comp}}}$ (16)
Here, $\mu_{\text{comp}}$ and $\sigma_{\text{comp}}$ are the corpus-wide mean and standard deviation of each component. This normalization transforms each sub-score to a distribution with mean 0 and standard deviation 1, enabling fair comparison and integration between components.
The normalized sub-scores are integrated through weighted linear combination:
$\tilde{S}_{\text{orig}}(x)=\lambda_{1}\tilde{S}_{\text{struct}}(x)+\lambda_{2}\tilde{S}_{\text{claim}}(x)+\lambda_{3}\tilde{S}_{\text{evid}}(x)+\lambda_{4}\tilde{S}_{\text{move}}(x)$ (17)
Here, $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}$ are the weights for each component, satisfying $\sum_{i=1}^{4}\lambda_{i}=1$. This study used $\lambda_{1}=\lambda_{2}=\lambda_{3}=\lambda_{4}=0.25$, a design that treats the four components equally.
The final score is calculated by multiplying the normalized originality score by the quality score:
$S_{\text{final}}(x)=\tilde{S}_{\text{orig}}(x)\times Q(x)$ (18)
The multiplicative design ensures that when $Q(x)\approx 0$, $S_{\text{final}}(x)\approx 0$ even if originality is high, preventing logically flawed essays from receiving high scores. On the other hand, when $Q(x)\approx 1$, the originality score is almost directly reflected in the final score. This multiplicative design positions quality as a prerequisite for originality.
Finally, Algorithm 1 shows the overall processing procedure of AROA. Through this series of processes, AROA enables evaluation of essays based on the originality of achieved thinking rather than surface fluency or author attribution, in an environment where AI-assisted essay writing has become commonplace.
Algorithm 1 AROA: Argument Rarity-based Originality Assessment
1:essay collection $\mathcal{D}=\{x_{i}\}_{i=1}^{N}$, weight parameters $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4},\alpha,\beta,\gamma,w_{\text{coh}},w_{\text{log}}$
2:Final score for each essay $\{S_{\text{final}}(x_{i})\}_{i=1}^{N}$
3:// Stage 1: Argument Structure Extraction
4:for each $x_{i}\in\mathcal{D}$ do
5:  $\mathcal{S}(x_{i})\leftarrow\text{LLM\_Extract}(x_{i})$ $\triangleright$ Extract 5 elements and relationships
6:  $\text{Depth}(x_{i}),n_{\text{matched}}(x_{i})\leftarrow\text{ComputeFeatures}(\mathcal{S}(x_{i}))$
7:end for
8:// Stage 2: Embedding
9:for each element $e\in\bigcup_{i}\mathcal{S}(x_{i})$ do
10:  $\mathbf{v}(e)\leftarrow\text{SentenceBERT}(e)$
11:end for
12:// Stage 3: Originality Sub-scores
[…]
16:for each $x_{i}\in\mathcal{D}$ do
17:  $\text{Coverage}(x_{i})\leftarrow n_{\text{matched}}(x_{i})/\max(n_{A}(x_{i}),1)$
18:  $S_{\text{move}}(x_{i})\leftarrow\alpha\cdot n_{\text{matched}}(x_{i})+\beta\cdot\text{Depth}(x_{i})+\gamma\cdot\text{Coverage}(x_{i})$
19:end for
20:// Stage 4: Quality Adjustment
21:for each $x_{i}\in\mathcal{D}$ do
22:  $Q_{\text{coherence}}(x_{i})\leftarrow\text{LLM\_Evaluate}(x_{i},\text{``coherence''})$
23:  $Q_{\text{logical}}(x_{i})\leftarrow\text{LLM\_Evaluate}(x_{i},\text{``logical''})$
24:  $Q(x_{i})\leftarrow w_{\text{coh}}\cdot Q_{\text{coherence}}(x_{i})+w_{\text{log}}\cdot Q_{\text{logical}}(x_{i})$
25:end for
26:// Stage 5: Final Score Computation
27:for each component $\text{comp}\in\{\text{struct},\text{claim},\text{evid},\text{move}\}$ do
28:  $\mu_{\text{comp}},\sigma_{\text{comp}}\leftarrow\text{Mean}(\{S_{\text{comp}}(x_{i})\}),\text{Std}(\{S_{\text{comp}}(x_{i})\})$
29:  $\tilde{S}_{\text{comp}}(x_{i})\leftarrow(S_{\text{comp}}(x_{i})-\mu_{\text{comp}})/\sigma_{\text{comp}}$ for all $i$
30:end for
31:for each $x_{i}\in\mathcal{D}$ do
32:  $\tilde{S}_{\text{orig}}(x_{i})\leftarrow\sum_{j=1}^{4}\lambda_{j}\tilde{S}_{j}(x_{i})$
33:  $S_{\text{final}}(x_{i})\leftarrow\tilde{S}_{\text{orig}}(x_{i})\times Q(x_{i})$
34:end for
35:return $\{S_{\text{final}}(x_{i})\}_{i=1}^{N}$
4 Experiments and Analysis
4.1 Dataset
This study used the AI Detection for Essays Dataset (AIDE) [6]. AIDE is a dataset constructed for AI-generated text detection in educational settings and contains both argumentative essays written by students and essays generated by AI. The essays concern two topics: Car-free cities and Electoral College, both of which are assignments requiring students to read given materials and argue their own position. There are two reasons for selecting this dataset. First, it contains numerous human essays on the same topic and same assignment, making it suitable for evaluating the proposed framework based on rarity within a corpus. Second, since it contains both human essays and AI essays, it enables verification of whether the proposed method can distinguish between original human argumentation and typical AI argumentation.
Table 2 shows the dataset statistics. Human essays were collected from student submissions. AI essays were those created following AIDE’s generation protocol. Specifically, 3-shot prompting was used, with three randomly selected human essays on the same topic as in-context examples, generating essays based on the same assignment instructions and reference materials. After generation, the following post-processing was applied: i) deduplication with a threshold of 85% cosine similarity using Universal Sentence Encoder embeddings, ii) manual removal of outputs that students would obviously not submit, iii) normalization of formatting and special characters, and iv) statistical matching of word count, sentence length, and function word distributions to human essays.
Table 2: Dataset statistics
Category Car-free cities Electoral College Total
Human essays 707 668 1,375
AI essays 500 500 1,000
Total 1,207 1,168 2,375
For the five-element extraction and quality evaluation, this study used GPT-4.1-mini [22]. GPT-4.1-mini has high instruction-following capability while having low API cost, making it suitable for processing large-scale datasets. Other parameters follow the settings described in Section 3.
4.2 Evaluation on Originality Score Distribution
This section analyzes the originality score distributions of human essays and AI essays to verify whether the proposed framework AROA can appropriately evaluate originality.
4.2.1 Overall Score Distribution
Figure 2 shows the distribution of final originality scores $S_{\text{final}}$ for human essays ($n=1,375$) and AI essays ($n=1,000$). Both distributions overlap around the mean, with the human essay mean at 0.015-0.015 and the AI essay mean at 0.001-0.001, which are nearly equivalent. What is noteworthy is the difference in distribution shapes. Human essays are concentrated in a relatively narrow range (standard deviation 0.345), while AI essays are dispersed over a wider range (standard deviation 0.534). This suggests that AI essay originality varies greatly depending on the prompting and sampling conditions.
Figure 2: Distribution of final originality scores ($S_{\text{final}}$) for human essays and AI essays. Dashed lines indicate the mean of each group.
Table 3 summarizes the statistical comparison. The final originality score $S_{\text{final}}$ showed no statistically significant difference between human essays and AI essays ($p=0.118$, Cohen’s d=0.032d=-0.032). This result has important implications. The proposed framework does not function as a mere AI detector but evaluates the originality of argumentation itself regardless of author attribution. Even human-written essays receive low scores if they remain at typical argumentation, while AI-generated essays receive high scores if they contain original argumentation.
Table 3: Statistical comparison between human and AI essays
Metric Human ($n=1,375$) AI ($n=1,000$) Cohen’s $d$ $p$-value
$S_{\text{final}}$ $015\pm 0.345 0.0$$01\pm 0.534 0.0$32-0.032 0.118
$S_{\text{struct}}$ $4.987\pm 0.962$ $4.797\pm 0.885$ 0.204 <0.001<0.001***
$S_{\text{claim}}$ $170\pm 0.076 0.0$$037\pm 0.025$ 2.239 <0.001<0.001***
$S_{\text{evid}}$ $092\pm 0.042 0.0$$048\pm 0.014$ 1.343 <0.001<0.001***
$S_{\text{move}}$ $5.093\pm 2.159$ $5.156\pm 2.340$ 0.028-0.028 0.821
$Q$ $0.609\pm 0.141$ $0.998\pm 0.021$ 3.600-3.600 <0.001<0.001***
4.2.2 Component-wise Analysis
Figure 3 shows the comparison between human essays and AI essays for each of the four originality components. Although overall scores are similar, both groups showed markedly different component profiles.
Figure 3: Comparison of four originality components between human and AI essays: Structural Rarity ($S_{\text{struct}}$), Claim Rarity ($S_{\text{claim}}$), Evidence Rarity ($S_{\text{evid}}$), and Cognitive Depth ($S_{\text{move}}$).
First, in claim rarity $S_{\text{claim}}$, human essays showed substantially higher scores (human: $0.170$, AI: $0.037$, $d=2.24$). As shown in the second panel of Figure 3, this difference is visually striking, with AI essay scores remaining at approximately one-fifth of human essays. This indicates higher diversity of perspectives and claims in human essays. AI-generated essays tend to converge on typical claim patterns contained in training data, resulting in a clustered distribution in semantic space.
A similar trend was observed in evidence rarity $S_{\text{evid}}$ (human: $0.092$, AI: $0.048$, $d=1.34$). Human essays used more diverse examples, statistics, and citations, while AI essays relied on commonly cited typical evidence. As the error bars indicate, human essays also show greater variation in evidence selection, suggesting the inclusion of unique evidence based on individual knowledge and experience.
On the other hand, no substantial differences were observed in structural rarity $S_{\text{struct}}$ (human: $4.987$, AI: $4.797$, $d=0.20$) and cognitive depth $S_{\text{move}}$ (human: $5.093$, AI: $5.156$, d=0.03d=-0.03). As shown in the first and fourth panels of Figure 3, the bar graphs for both groups are nearly the same height, and the error bars overlap. This indicates that AI can sufficiently reproduce the structural complexity in human argumentation (hierarchical structure of claim-reason-evidence and dialogic structure of counterargument-rebuttal).
4.3 Evaluation on Quality-Originality Relationship
This section analyzes the relationship between quality and originality and verifies the validity of the multi-component design of the proposed framework AROA.
4.3.1 Component Independence
Figure 4 shows the correlation matrix among originality components and quality score. First, structural rarity $S_{\text{struct}}$ has extremely low correlations with all other components (r=0.08r=0.080.130.13). This indicates that the atypicality of argument structure is a dimension independent of claim/evidence content and depth of reasoning. In other words, essays with structurally rare argumentation patterns do not necessarily contain rare claims or evidence. This result supports the design philosophy of this framework that how arguments are made and what is argued are different evaluation axes.
Next, cognitive depth $S_{\text{move}}$ also has low correlations with other components (r=0.06r=0.060.130.13). The dialogic structure of counterargument-rebuttal and hierarchical depth of reasoning vary independently of claim/evidence rarity. Essays with deep reasoning structures do not necessarily make rare claims, and vice versa.
On the other hand, a moderate positive correlation ($r=0.51$) was observed between claim rarity $S_{\text{claim}}$ and evidence rarity $S_{\text{evid}}$. This is an intuitively understandable result. Authors who make claims from original perspectives tend to also explore and present unique evidence to support those claims. Both components share the common property of semantic rarity while evaluating different argument elements (claims vs. evidence), so they do not completely coincide.
These results support the validity of multifaceted evaluation using four components. If all components showed high correlations, the significance of using multiple components would diminish. However, in practice, each component captures different aspects of originality, enabling multifaceted evaluation that is not visible with a single indicator.
Figure 4: Correlation matrix of originality components and quality score.
4.3.2 Quality Does Not Guarantee Originality
The most striking finding in the correlation matrix is the strong negative correlation between quality score $Q$ and semantic rarity components. The correlation between $Q$ and $S_{\text{claim}}$ is r=0.67r=-0.67, and between $Q$ and $S_{\text{evid}}$ is r=0.50r=-0.50, both showing statistically significant strong negative correlations. This result means that higher-quality essays tend to have lower rarity of claims and evidence.
Behind this negative correlation lies a trade-off between quality and rarity. To write logically coherent, high-quality text, it is easier to use widely accepted claim patterns and commonly cited evidence. Using original claims and rare evidence makes it difficult to maintain logical coherence, and quality scores tend to decrease.
On the other hand, correlations between $Q$ and structural components ($S_{\text{struct}}$: r=0.10r=-0.10, $S_{\text{move}}$: r=0.08r=-0.08) are extremely weak. This indicates that the complexity and atypicality of argument structure do not significantly affect quality. Even essays with complex structures can maintain logical coherence.
Figure 5 shows the human-AI comparison of quality score $Q$ and its components ($Q_{\text{coherence}}$, $Q_{\text{logical}}$). AI essays substantially exceeded human essays in all quality metrics. In particular, AI essay quality scores are nearly perfect ($Q=0.998$), achieving the highest levels in both coherence ($Q_{\text{coherence}}\approx 1.0$) and logical validity ($Q_{\text{logical}}\approx 1.0$). On the other hand, human essays remain at $Q=0.609$, with room for improvement particularly in logical validity ($Q_{\text{logical}}\approx 0.50$).
Figure 5: Comparison of quality scores between human and AI essays.
These results clearly demonstrate that text quality does not guarantee originality. AI essays generate fluent and logically coherent text, but their high quality is achieved by relying on typical claim patterns and common evidence. The strong negative correlation between $Q$ and $S_{\text{claim}}$$S_{\text{evid}}$ shown in the correlation matrix quantitatively supports this trade-off. The reason AI essays that would receive high scores in conventional writing assessment remain at originality scores equivalent to human essays in this framework is that it appropriately captures this quality-rarity trade-off. The proposed framework appropriately models the relationship between quality and originality through the design that multiplies quality score $Q$ into the final score (Equation 18). In AI essays, the low originality components are partially compensated by high $Q$, while in human essays, high originality components are adjusted by the variation in $Q$. As a result, the final score distributions of both groups become similar, reflecting that quality and originality are independent evaluation axes.
4.3.3 Illustrative Examples
To concretely illustrate the relationship between quality and originality, representative essays were extracted from the AIDE dataset and their scores and content were compared. Table 4 shows the scores of high-originality and low-originality essays.
Table 4: Comparison of high-originality and low-originality essays
Essay $S_{\text{final}}$ $S_{\text{struct}}$ $S_{\text{claim}}$ $S_{\text{evid}}$ $S_{\text{move}}$ $Q$
High-originality 1.410 7.444 0.294 0.135 10.0 0.75
Low-originality 1.230-1.230 4.100 0.038 0.046 3.0 1.00
The high-originality essay ($S_{\text{final}}=1.41$) argues from multiple perspectives on the topic of Car-free cities. The following is an excerpt:
“having a society like this is also very stress relieving in a sense. you no longer have the every day stress of being on a highway or constantly alert for cars […] the amount of exercise is also increased here because you are walking or biking everywhere”
“Now with all this being said there obviously are still downsides to cutting out gas producing cars. weather being a main one. suppose it rains one day, how do you get to work? […] if you work out of town how do you get to your job?”
There are three reasons why this essay received a high score. First, it presents original perspectives such as stress reduction and increased exercise, not just environmental issues, leading to a high claim rarity of $S_{\text{claim}}=0.294$. Second, by raising counterarguments (weather, long-distance commuting, transporting work tools) and responding to them, it achieves a deep cognitive depth of $S_{\text{move}}=10.0$. Third, although the quality score is not perfect at $Q=0.75$, the original content compensates for it.
On the other hand, the low-originality essay ($S_{\text{final}}=-1.23$) takes a typical approach to the same topic:
“The actual amount of people driving and getting their license has decreased over the past couple years and this is due to two reasons, not driving helps the environment and it helps the community.”
“There has been large amounts of information drawn from certain ‘Car-Free’ experiments around the world. It has shown that the limited usage of cars has improved the environment by reducing smog in Paris and has also helped communities like Vauban increase the happiness of its citizens.”
This essay is logically coherent and achieves $Q=1.00$, but has the lowest originality score. There are three reasons for this. First, claims such as “good for the environment” and “good for the community” are stereotypical patterns frequently seen in other essays addressing the same topic, leading to a low claim rarity of $S_{\text{claim}}=0.038$. Second, throughout the essay, it takes the form of listing citations from materials, lacking original analysis or specific examples ($S_{\text{evid}}=0.046$). Third, it does not consider any counterarguments to its claims and remains at a simple “reason 1, reason 2” structure, resulting in the minimum cognitive depth of $S_{\text{move}}=3.0$.
These concrete examples clearly show what the proposed framework evaluates. High originality scores arise from original perspectives, diverse evidence, and deep argumentation structures that anticipate and respond to counterarguments. On the other hand, essays that are logically coherent but rely on typical claims and common evidence receive low scores. This succinctly demonstrates the difference between conventional writing assessment (quality-focused) and this framework (originality-focused).
4.4 Evaluation on Robustness to Sample Size
This section verifies the stability of scores against reference corpus size. Since the proposed method is based on density estimation, the reference corpus size may affect score reliability. Using the entire human essay dataset as a reference, subsampling experiments were conducted. For each sample size $n\in\{30,50,100,200,300,500,700,1000\}$, the following procedure was repeated 5 times with different random seeds: i) randomly sample $n$ essays from the full dataset, ii) recalculate structural rarity scores $S_{\text{struct}}$ using only the subsample, and iii) compute Spearman’s rank correlation coefficient $\rho$ between rankings based on the subsample and rankings based on the full data. $S_{\text{struct}}$, which depends on density estimation, is the component most sensitive to sample size, and if this is stable, other components can be expected to be similarly stable.
Figure 6 shows the relationship between sample size and rank correlation. Rank correlation improves monotonically with increasing sample size, exceeding the $\rho=0.80$ threshold around $n=300$ and the $\rho=0.90$ threshold around $n=700$. As the error bars indicate, variation between trials also decreases as sample size increases. On the other hand, at typical classroom scales (n=30n=305050), rank correlation remains at approximately 0.50. These results indicate that when actually applying the proposed method, the reference corpus size needs to be considered.
Figure 6: Rank correlation stability against sample size. The plot shows Spearman’s rank correlation ($\rho$) between subset-based and full-data-based rankings at each sample size. Error bars indicate standard deviation across 5 trials.
4.5 Cost Analysis
This section analyzes computational cost and API cost for practical deployment of the proposed framework. The experiments were conducted on a macOS environment equipped with an Apple M1 chip (8-core CPU, 7-core GPU) and 16GB RAM, using 10 parallel workers.
Table 5 shows the processing time breakdown per essay. Approximately 90% of processing time is occupied by five-element extraction through LLM API calls. This reflects that the process of analyzing the entire essay and structuring claims, reasons, evidence, counterarguments, and rebuttals is the most computationally intensive. On the other hand, text embedding and density estimation are negligible as local processing, making it an efficient design from an API cost perspective as well.
Table 5: Processing time breakdown per essay
Component Time Note
Argument extraction 1.35 sec GPT-4.1-mini API
Text embedding 0.04 sec Local Sentence-BERT
Density estimation <<0.01 sec NumPy/scikit-learn
Quality evaluation 0.13 sec GPT-4.1-mini API
Total 1.52 sec
Table 6 shows measured values from this experiment and estimated costs assuming application at various scales. The measured time for processing 1,375 human essays was 34.8 minutes with an API cost of \$3.36, approximately \$0.0024 (about 0.24 cents) per essay. From classroom scale (30 essays) to institution scale (10,000 essays), processing time and cost scale linearly, enabling operation at realistic costs even for large-scale applications.
Table 6: Scalability analysis: processing time and API cost
Scale Essays Time API Cost
Single class 30 0.8 min \$0.07
Multiple classes 100 2.5 min \$0.24
Course level 500 13 min \$1.20
Department level 1,375 35 min \$3.36 (measured)
Large-scale 5,000 2.1 hours \$12.00
Institution level 10,000 4.2 hours \$24.00
In comparison with human evaluation, the efficiency of the proposed method is remarkable. When human evaluators manually evaluate essay originality, it typically takes 5–10 minutes per essay. In contrast, the proposed method can process in approximately 1.5 seconds, which is 200–400 times faster. In terms of cost, approximately \$0.0024 per essay compares to \$1–5 for human evaluation depending on evaluator expertise and evaluation depth, making it approximately 400–2,000 times cheaper. This efficiency enables large-scale and frequent originality assessment that was previously difficult to implement.
4.6 Evaluation on Transferability Across LLMs
This section conducts transferability analysis using three models, GPT-4.1-mini, Gemini 2.5 Flash Lite, and Claude 3 Haiku, to verify whether the proposed framework depends on the specific LLM used for argument structure extraction.
Table 7 shows the cross-model correlations of final originality score $S_{\text{final}}$. The average Pearson correlation across all model pairs was $r=0.465$, confirming moderate agreement in originality evaluation. GPT showed relatively high agreement with both Gemini and Claude, while the correlation between Gemini and Claude was the lowest ($r=0.402$), suggesting that these two models emphasize different aspects of argumentative originality.
Table 7: Cross-model correlation of $S_{\text{final}}$
Model pair Pearson $r$ Spearman $\rho$
GPT–Gemini 0.505 0.501
GPT–Claude 0.489 0.496
Gemini–Claude 0.402 0.416
Average 0.465 0.471
Table 8 shows the cross-model agreement for each component score. Notable differences in transferability were observed depending on the component. Evidence rarity $S_{\text{evid}}$ showed the highest stability (average $r=0.616$), confirming that evidence identification and evaluation are relatively consistent across different LLMs. On the other hand, structural rarity $S_{\text{struct}}$ showed the lowest agreement (average $r=0.232$), revealing that argument structure recognition differs significantly by LLM. Quality evaluation $Q$ also showed low transferability (average $r=0.249$). In particular, Claude assigned nearly uniform quality scores to almost all essays, suggesting that quality evaluation criteria fundamentally differ between models.
Table 8: Cross-model correlation by component (Pearson $r$)
Component GPT–Gemini GPT–Claude Gemini–Claude Average Stability
$S_{\text{evid}}$ 0.658 0.626 0.563 0.616 High
$S_{\text{claim}}$ 0.462 0.459 0.532 0.484 Medium
$S_{\text{move}}$ 0.475 0.460 0.352 0.429 Medium
$Q$ 0.483 0.114 0.148 0.249 Low
$S_{\text{struct}}$ 0.294 0.268 0.133 0.232 Low
Figure 7 shows the contribution (correlation coefficient) of each component to $S_{\text{final}}$ for each model. Characteristic evaluation tendencies were observed for each model. GPT-4.1-mini shows nearly equal contributions across all four components (r=0.54r=0.540.560.56), performing balanced evaluation. Gemini 2.5 Flash Lite emphasizes structural rarity $S_{\text{struct}}$ the most ($r=0.55$) while showing relatively low sensitivity to evidence rarity $S_{\text{evid}}$ ($r=0.43$). Claude 3 Haiku places the greatest weight on claim rarity $S_{\text{claim}}$ ($r=0.62$) and evidence rarity $S_{\text{evid}}$ ($r=0.58$), showing a pronounced tendency to emphasize semantic rarity.
Figure 7: Component impact on final score by model.
These results indicate that the proposed framework captures the core constructs of argumentative originality regardless of the LLM adopted, while revealing that component stability differs systematically. Evidence-related scores transfer well across models, while structure and quality evaluations are more model-dependent. The differences in evaluation characteristics of each model and their practical implications are discussed in Section 5.
5 Discussion
5.1 Key Findings
This study proposed AROA, a framework for automatically evaluating argumentative originality in essays, and verified its effectiveness through experiments using 1,375 human essays and 1,000 AI essays. This section discusses the key findings from the experimental results and their implications.
The most important finding is that text quality and originality are independent evaluation axes, and a trade-off exists between them. In the correlation analysis, strong-to-moderate negative correlations were observed: r=0.67r=-0.67 between quality score $Q$ and claim rarity $S_{\text{claim}}$, and r=0.50r=-0.50 between $Q$ and evidence rarity $S_{\text{evid}}$. This result suggests that to write logically coherent, high-quality text, it is easier to use widely accepted claim patterns and common evidence. Using original claims and rare evidence makes justification and maintenance of logical coherence difficult, resulting in lower quality scores. This finding highlights the problem of conventional writing assessment being biased toward quality. Fluent and logically coherent text does not necessarily reflect original thinking; rather, it may indicate dependence on typical patterns.
The comparison between human essays and AI essays vividly demonstrates this quality-originality trade-off. While AI essays achieved near-perfect quality scores ($Q=0.998$), their claim rarity remained at approximately one-fifth of human essays (AI: 0.037, human: 0.170). Interestingly, no significant differences were observed between the two groups in structural rarity $S_{\text{struct}}$ and cognitive depth $S_{\text{move}}$. This means that contemporary LLMs can sufficiently reproduce human argumentation structures—the hierarchical structure of claim-reason-evidence and the dialogic structure of counterargument-rebuttal. However, in the semantic dimension of what to claim and what evidence to choose, AI tends to converge on typical patterns contained in training data, finding it difficult to generate unique evidence based on diverse perspectives and personal experiences that humans possess. This finding provides a perspective different from AI-generated text detection. The proposed framework AROA does not determine whether the author is AI or human but evaluates the originality of argumentation itself. In fact, no statistically significant difference was observed in the final score $S_{\text{final}}$ between the two groups ($p=0.118$). This means that even human-written essays receive low scores if they remain at typical argumentation, while AI-generated essays receive high scores if they contain original argumentation.
The validity of multifaceted evaluation using four components was also confirmed. In the correlation analysis, structural rarity $S_{\text{struct}}$ and cognitive depth $S_{\text{move}}$ showed extremely low correlations with other components (r=0.06r=0.060.130.13), indicating that each captures different aspects of originality. How arguments are made (structure) and what is argued (semantic content) are independent evaluation axes, enabling multifaceted originality evaluation that cannot be captured by a single indicator. On the other hand, a moderate positive correlation ($r=0.51$) was observed between claim rarity $S_{\text{claim}}$ and evidence rarity $S_{\text{evid}}$. This indicates that authors who make claims from original perspectives tend to also explore and present unique evidence to support those claims, confirming that while sharing the common property of semantic rarity, they evaluate different argument elements.
From a practical perspective, this framework has sufficient efficiency and reliability for deployment in educational settings. Processing costs of approximately 1.5 seconds and approximately \$0.0024 per essay make large-scale evaluation realistic. Robustness analysis showed that rank correlation $\rho\geq 0.80$ is achieved when the reference corpus size is 300 or more, yielding reliable results for undergraduate-level courses and evaluations pooling multiple classes. This finding provides a rationale for large-scale course designs. The RaaS implementation at Ritsumeikan University, which accommodates up to 1,000 students per class as mentioned in Section 1, is well-suited for density-based originality assessment. Larger cohorts provide richer reference corpora, enabling more reliable rarity estimation. In this sense, the seemingly challenging scale of large enrollment courses becomes an advantage rather than a limitation when combined with AI-assisted originality assessment. On the other hand, at typical classroom scales (30–50 essays), rank correlation remains at approximately 0.50, so for small-scale classes, it is recommended to use the framework for formative feedback rather than summative ranking.
Transferability analysis confirmed that the proposed framework does not heavily depend on a specific LLM. Among the three models—GPT-4.1-mini, Gemini 2.5 Flash Lite, and Claude 3 Haiku—the average correlation of final scores was $r=0.465$, showing moderate agreement. By component, evidence rarity $S_{\text{evid}}$ showed the highest stability (average $r=0.616$), while structural rarity $S_{\text{struct}}$ and quality evaluation $Q$ showed high model dependency (average r=0.23r=0.230.250.25). Each model showed characteristic evaluation tendencies: GPT performed balanced evaluation, Gemini emphasized structure, and Claude emphasized semantic rarity. This result suggests the possibility of more robust evaluation through an ensemble of multiple models.
These findings suggest a paradigm shift in writing assessment in the AI era. Now that LLMs can easily generate high-quality text, assessment based solely on quality is losing its significance. What is important is whether the author has original perspectives, explores diverse evidence, and has gone through a deep thinking process that anticipates and responds to counterarguments. However, it is not easy for human evaluators to consistently evaluate such originality. When grading numerous essays, evaluators are susceptible to order effects and fatigue, where impressions from earlier essays influence subsequent evaluations. Also, evaluators tend to give higher scores to claims that match their prior knowledge and expectations, while finding it difficult to appropriately recognize the value of unfamiliar perspectives or rare evidence. Furthermore, considering relative rarity within the entire corpus when evaluating individual essays is cognitively difficult. The proposed framework AROA complements these limitations of human evaluation. Through density-based rarity evaluation, it can objectively and consistently quantify how rare each essay’s claims and evidence are within the entire corpus. It can evaluate all essays according to the same criteria without being affected by evaluator fatigue or order effects. This provides a foundation for educators to more fairly and efficiently evaluate and develop students’ critical thinking skills while referring to the proposed framework’s scores.
5.2 Limitations
This study has several limitations. First, the evaluation is based on a single dataset and is limited to two topics. Generalizability to different topics, different languages, and different grade levels requires future verification. In particular, applicability to genres other than argumentative essays, such as specialized academic fields and creative writing, remains unverified.
Second, the proposed framework has not been validated for correlation with originality evaluation by human experts. This study theoretically discussed that density-based rarity is a valid operational definition of originality and demonstrated its validity through concrete examples, but future empirical research is needed to determine how well this framework’s scores align with essays that educators and researchers intuitively judge as original.
Third, since argument structure extraction and quality evaluation depend on LLMs, LLM biases and limitations may affect scores. Although transferability analysis confirmed moderate cross-model agreement, high model dependency was shown particularly in structural rarity and quality evaluation. Also, since LLM capabilities are rapidly evolving, the results from models used in this study may not directly apply to future models.
Fourth, there is dependency on reference corpus size. Robustness analysis showed that reliable evaluation is possible with 300 or more essays, but there are constraints on application in small-scale classes. Further consideration is also needed regarding how the composition of the reference corpus (students’ grade levels, assignment difficulty, etc.) may affect scores.
Future work should address these limitations and consider the following directions: i) verification of generalizability using diverse datasets, ii) empirical validation through correlation analysis with human evaluation, iii) application to formative assessment through development of feedback generation functionality, and iv) extension to longitudinal studies tracking the development of originality over time. Through these extensions, the proposed framework is expected to contribute further to the evaluation and development of critical thinking skills in educational settings.
6 Conclusion
As LLMs have become capable of effortlessly generating high-quality text, traditional quality-focused writing assessment is losing its significance. This study proposed AROA, a framework for automatically evaluating argumentative originality in essays. AROA operationally defines originality as rarity within a reference corpus and evaluates it through four complementary components: structural rarity, claim rarity, evidence rarity, and cognitive depth. The framework quantifies the rarity of each component using density estimation and adjusts them with quality scores, thereby treating quality and originality as independent evaluation axes.
Experiments using 1,375 human essays and 1,000 AI essays verified the effectiveness of the proposed framework. A strong negative correlation (r=0.67r=-0.67) was observed between quality and claim rarity, revealing a quality-originality trade-off where higher-quality texts tend to rely on typical claim patterns. Furthermore, while AI essays achieved comparable levels of structural complexity to human essays, their claim rarity was approximately one-fifth that of humans, indicating that LLMs can reproduce the form of argumentation but have limitations in the originality of content.
Future research directions include verification of generalizability across diverse datasets, empirical validation through correlation analysis with human evaluation, and application to formative assessment through development of feedback generation functionality. This study provides a foundation for a paradigm shift in writing assessment in the AI era, from quality to originality.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgment
This work was supported by a research grant from the Ritsumeikan Global Innovation Research Organization (R-GIRO), Ritsumeikan University. The authors also acknowledge the RaaS (Ritsumeikan AI-powered Assessment Solution) system developed and deployed at Ritsumeikan University, which provided practical insights into large-scale AI-assisted assessment (https://raas.ritsumei.ac.jp/).
Data availability
The AIDE dataset used in this study is publicly available at Kaggle (https://www.kaggle.com/datasets/lburleigh/tla-lab-ai-detection-for-essays-aide-dataset). The code for AROA will be made available upon request.
Declaration of Generative AI Use
During the preparation of this manuscript, the authors used generative AI tools for language editing and organization. All content was reviewed and verified by the authors, who take full responsibility for the final manuscript.